OpenStudy (anonymous):

simp. in Rad. form. Sqrt 28

OpenStudy (saifoo.khan):

\[2 \sqrt7\]

OpenStudy (anonymous):

i thought it was 4\[\sqrt{7}\]

OpenStudy (anonymous):

so its 2sqrt7?

OpenStudy (anonymous):

thx

OpenStudy (saifoo.khan):

No, there will be a \[2 \sqrt7\]

OpenStudy (saifoo.khan):

PLease hit good answer.

OpenStudy (anonymous):

we can rewrite it as...sqrt(28)=sqrt(2*2*7)=(sqrt(2)^2*sqrt7)=2sqrt7

OpenStudy (saifoo.khan):

Correct.

OpenStudy (anonymous):

as we know that square root of a square will cancel out each other with result 1.

OpenStudy (anonymous):

ok thx tht explains a lot. thx to the both of you.

OpenStudy (anonymous):

wouldnt sqrt384 be 2sqrt7?

OpenStudy (anonymous):

just seeing if i have the hang of this

OpenStudy (anonymous):

Here, we can write the given expression as... sqrt(384)=sqrt(2*2*2*2*2*2*2*3)=sqrt(2^2*2^2*2^2*2*3)=2*2*2*sqrt(6)=8sqrt(6) so..a big no for your answer

OpenStudy (anonymous):

OK i see where i went wrong... i wrote 14 not 12. thx

OpenStudy (anonymous):

how would you do 6*sqrt72

OpenStudy (anonymous):

plz?

OpenStudy (anonymous):

here we can rewrite our equation as.. 6sqrt72=6*sqrt(2*2*2*3*3)=6*sqrt(2^2*3^2*2)=6*2*3sqrt(2) 36sqrt2...

OpenStudy (anonymous):

howd you do tht. Explain.

OpenStudy (anonymous):

here i just expend the sqrt term first...sqrt72 i started doing division by taking the least number 2 till it is not divisble more...and then 3 and then 4 and so on... i got sqrt72=6sqrt2 now, i simpli multiplied it to 6 so..6*6sqrt2=36sqrt2

OpenStudy (anonymous):

sqrt72=sqrt(2*2*2*3*3)=sqrt(2^2*3^2*2)=2*3sqrt(2)=6sqrt2

OpenStudy (anonymous):

72/2=36 36/2=18 18/2=9 9/3=3 3/3=1 so...count the number of 2's and 3's needed to divide the number 72 here...and multiply then together to get the factor of the number 72

OpenStudy (anonymous):

see i dontget that i understand how to break it down but not how to get the # to multiply it by would it be 2*2*2*3*3?

OpenStudy (anonymous):

yes...exactly..!:)

OpenStudy (anonymous):

you got it.......great!

OpenStudy (anonymous):

so tht would be 72? thn what would you do ?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

as we know that the square root of square of any number cancel out each other....like...sqrt(2^2) as we know that we can rewrite the sqrt function as..sqrt(2^2)=(2^2)^(1/2) so, it will become...(2)^(2*1/2).. here 2 in power cancel out each other ...we have left only 2 here...so...we need to sort out all the term from the square root here in the same way..we will get...sqrt72=sqrt(2*2*2*3*3)=sqrt(2^2*3^2*2)=(sqrt(2^2)*sqrt(3^2)*sqrt(2)=2*3sqrt2=6sqrt2 HOPE THIS WILL HELP YOU

OpenStudy (anonymous):

(sqrt2)^2=sqrt(2^2)=2..let me know if you need more help on that?